OpenStudy (anonymous):

A scientist counts 50 bacteria present in a culture and finds that the number of the bacteria in the culture doubles each hour. What is the function that models the number of bacteria after x hours?

OpenStudy (anonymous):

50*(2^x) i think cause if its the first hour it would be 50*2, then (50*2)*2, then ((50*2)*2)*2 so basically 50*2^3 and so on... :)

OpenStudy (whpalmer4):

Yes, you can write a doubling equation like this as \[P(t) = P_0*2^t\]where \[P(0) = P_0\] Sometimes you'll see it written as \[P(t) = P_0 e^{\lambda t}\]where \[\lambda = \frac{\ln 2}{t_2} \]and \(t_2\) is the time it takes to double.